# Toeplitz Operators Defined between Köthe Spaces

### v1

@@ 3. Operators Defined by A Lower Triangular Toeplitz Matrix
If $\widehat{T}_{\theta}$ is compact, then $\widehat{T}_{\theta}$ is continuous and $\theta\in\Lambda_{1}(\beta)$ from Proposition 3.1. Lemma 2.1 says that there exists a $m\in N$ such that for all $k\in\mathbb{N}$
This gives us that for every $k\in\mathbb{N}$ there exists a $C_{1}>0$ so that
$\|\widehat{T}_{\theta}e_{n}\|_{k}=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-{1\over k}\beta_{j}}\leq C_{1}\|e_{n}\|_{m}=C_{1}a_{n,m}$
By assuming $\theta_{0}\neq 0$, we can write
$e^{-{1\over k}\beta_{n}}\leq{C\over|\theta_{0}|}a_{n,m}$
for all $k,n\in\mathbb{N}$. This says that 3.2 is satisfied. For the converse, assume that $\theta\in\Lambda_{1}(\beta)$ and 3.2 holds. By the same calculation of the proof of Theorem 3.2, we can write that there exists $m\in\mathbb{N}$ such that for all $k\in\mathbb{N}$
@@
$\quad\quad\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad\quad e^{-\frac{1}{k}\beta_{n}}\leq Ca_{n,S(k)}\quad\quad\quad\forall n\in\mathbb{N}.$ (3.3)
@@
is S-tame.
@@
Let $\theta\in\Lambda_{1}(\beta)$. By using 3.3, for every $k\in\mathbb{N}$, there exist $C_{1},C_{2}>0$ such that
$\begin{split}\|\widehat{T}_{\theta}e_{n}\|_{k}&\leq\|\widehat{T}_{\theta}e_{n}\|_{2k}=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{2k}\beta_{j}}=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{k}\beta_{j}}e^{-\frac{1}{k}\beta_{j}}\\ &\leq\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{k}\beta_{j}}\bigg{)}e^{-\frac{1}{k}\beta_{n}}\leq C_{1}\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{k}\beta_{j}}\bigg{)}a_{n,S(k)}\\ &\leq C_{1}\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{k}\beta_{j-n+1}}\bigg{)}\|e_{n}\|_{S(k)}\leq C_{1}\|\theta\|_{k}\|e_{n}\|_{S(k)}\leq C_{2}\|e_{n}\|_{S(k)}.\end{split}$
@@
$\|\widehat{T}_{\theta}e_{n}\|_{k}\leq C\|e_{n}\|_{S(k)},$
that is, the family $\mathcal{\widehat{A}}_{1}$ is S-tame. ∎
Now we turn our attention to the power series space of infinite type. For the converse of Proposition 3.1, we need the following growth condition on the sequence $\beta$:
$\exists M\in\mathbb{N}\quad\forall t,s>t\quad\quad\beta_{s}\leq M(\beta_{s-t}+\beta_{t}).$ (3.4)
when $K(b_{n,k})$ is power series space of infinite type $\Lambda_{\infty}(\beta)$.
@@
Let $\beta=\left(\beta_{n}\right)_{n\in\mathbb{N}}$ be a non-negative increasing sequence with $\lim_{n\rightarrow\infty}\beta_{n}=\infty$. Assume that $\beta$ satisfies the condition in 3.4. $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{\infty}(\beta)$ is continuous if and only if $\theta\in\Lambda_{\infty}(\beta)$ and the following condition holds:
$\quad\quad\quad\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad\quad e^{k\beta_{n}}\leq Ca_{n,m}\quad\quad\forall n\geq k.$ (3.5)
@@
The sufficiency part comes from Proposition 3.1. For the necessary part, assume $\theta\in\Lambda_{\infty}(\beta)$ and 3.5 is satisfied. Then, for every $k\in\mathbb{N}$ there exist $m\in\mathbb{N}$, $C_{1},C_{2}>0$ such that
$\begin{split}\|\widehat{T}_{\theta}e_{n}\|_{k}&=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{k\beta_{j}}\leq\sum^{\infty}_{j=n}|\theta_{j-n}|e^{Mk(\beta_{j-n}+\beta_{n})}\\ &\leq\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{Mk\beta_{j-n+1}}\bigg{)}e^{Mk\beta_{n}}\\ &\leq C_{1}\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{Mk\beta_{j-n+1}}\bigg{)}a_{n,m}\\ &=C_{1}\|\theta\|_{Mk}\|e_{n}\|_{m}\\ &\leq C_{2}\|e_{n}\|_{m}\end{split}$
Therefore, for all $k\in\mathbb{N}$ there exist $m\in\mathbb{N}$ such that
$\sup_{n\geq k}\frac{\|\widehat{T}_{\theta}e_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty$
and this implies that
@@
that is, $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{\infty}(\beta)$ is continuous. ∎
Remark 3.6.
By using Binomial expansion, one can show that the sequence $\beta=(n^{\frac{1}{d}})_{n\in\mathbb{N}}$ satisfies the condition 3.4 for any $d\in\mathbb{N}$. Every sequence $\beta$ satisfying 3.4 is stable, that is, $\sup_{n\in\mathbb{N}}\frac{b_{2n}}{b_{n}}<\infty$. However the converse is not true since the sequence $\beta=(n^{2})_{n\in\mathbb{N}}$ is stable but it does not satisfy the condition 3.4. So that the condition in 3.4 is stronger than being stable.
@@
Let $\beta=\left(\beta_{n}\right)_{n\in\mathbb{N}}$ be a non-negative increasing sequence with $\lim_{n\rightarrow\infty}\beta_{n}=\infty$. Assume that $\Lambda_{\infty}(\beta)$ is nuclear and $\beta$ satisfies the condition in 3.4. $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{\infty}(\beta)$ is compact if and only if $\theta\in\Lambda_{\infty}(\beta)$ and the following condition holds:
$\quad\quad\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad\quad e^{k\beta_{n}}\leq Ca_{n,m}\quad\quad\forall n\geq k.$ (3.6)
@@
If $\widehat{T}_{\theta}$ is compact, then $\theta\in\Lambda_{1}(\beta)$ from Proposition 3.1. Lemma 2.1 says that there exists a $m\in N$ such that for all $k\in\mathbb{N}$
This gives us that there exists a $C_{1}>0$ such that for every $k\in\mathbb{N}$
$\|\widehat{T}_{\theta}e_{n}\|=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{k\beta_{j}}\leq C_{1}\|e_{n}\|_{m}=C_{1}a_{n,m}$
for every $n\geq k$. By assuming $\theta_{0}\neq 0$, we can write that for every $k\in\mathbb{N}$ we can find a $C>0$ such that
$e^{k\beta_{n}}\leq Ca_{n,m}$
for every $n\geq k$. This says that the equation 3.6 is satisfied. For the converse, assume that $\theta\in\Lambda_{\infty}(\beta)$ and 3.6 holds. By the same calculation of the proof of Theorem 3.5, we can write that there exists $m\in\mathbb{N}$ such that for all $k\in\mathbb{N}$
@@
Let $\beta=\left(\beta_{n}\right)_{n\in\mathbb{N}}$ be a non-negative increasing sequence with $\lim_{n\rightarrow\infty}\beta_{n}=\infty$. Assume that $\beta$ satisfies the condition in 3.4 and there exists a non-decreasing function $S:\mathbb{N}\to\mathbb{N}$ such that
$\quad\quad\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad\quad\quad e^{k\beta_{n}}\leq Ca_{n,S(k)}\quad\quad\quad\forall n\in\mathbb{N}.$ (3.7)
@@
is $\tilde{S}$-tame, where $\tilde{S}(k)=MS(k)$ for every $k\in\mathbb{N}$ and M comes from the condition 3.4.
@@
Let $\theta\in\Lambda_{\infty}(\beta)$. By using 3.7, we can write that for all $k\in\mathbb{N}$ there exist $C_{1},C_{2}>0$ so that for every $n\in\mathbb{N}$ the following inequalities hold:
$\begin{split}\|\widehat{T}_{\theta}e_{n}\|_{k}&=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{k\beta_{j}}\leq\sum^{\infty}_{j=n}|\theta_{j-n}|e^{Mk(\beta_{j-n}+\beta_{n})}\\ &\leq\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{Mk\beta_{j-n+1}}\bigg{)}e^{Mk\beta_{n}}\\ &\leq C_{1}\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n+1}|e^{Mk\beta_{j-n+1}}\bigg{)}a_{n,MS(k)}\\ &=C_{1}\|\theta\|_{kM}\|e_{n}\|_{MS(k)}\\ &\leq C_{2}\|e_{n}\|_{MS(K)}\end{split}$
@@
$\forall k\quad\exists C>0\quad\quad\|\widehat{T}_{\theta}e_{n}\|_{k}\leq C\|e_{n}\|_{\tilde{S}(k)}$
@@ 4. Operators Defined by an Upper Triangular Toeplitz Matrix
[…]
@@
In this section, we examine the properties of operators defined by an upper triangular Toeplitz matrix between Köthe spaces. In [8], you can find the same idea for Proposition 4.1 and a generalized form of Theorem 4.2 for $G_{\infty}$-spaces arguing as in Theorem 2.2
@@
Let $K(a_{n,k})$ and $K(b_{n,k})$ be Köthe spaces. If $\widecheck{T}_{\theta}:K(a_{n,k})\to K(b_{n,k})$ is continuous, then $\theta\in(K(a_{n,k}))^{\prime}$ and the following holds:
$\quad\quad\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad\quad\quad b_{n,k}\leq Ca_{n,m}\quad\quad\quad\quad\quad\forall n\in\mathbb{N}.$
We can say that $K(a_{n,k})\subseteq K(b_{n,k})$ in case (ii) is satisfied.
@@
Since $(b_{n,k})$ is a Köthe matrix, there exists a $k_{0}\in\mathbb{N}$ satisfying $b_{1,k_{0}}\neq 0$. Let take $C_{2}={C_{1}\over b_{1,k_{0}}}$. Then we can write
@@
and then this says that $\theta\in(K(a_{n,k}))^{\prime}$. We can assume that $\theta_{0}\neq 0$. Then, by choosing $C_{3}=\frac{C_{1}}{|\theta_{0}|}$ we can write
$b_{n,k}\leq C_{3}a_{n,m}\quad\quad\quad\forall n\in\mathbb{N}.$
@@
Let $\Lambda_{\infty}(\alpha)$ be a power series space of infinite type and $K(b_{n,k})$ be a nuclear Köthe space. The operator $\widecheck{T}_{\theta}:\Lambda_{\infty}(\alpha)\to K(b_{n,k})$ is continuous if and only if $\theta\in(\Lambda_{\infty}(\alpha))^{\prime}$ and the following condition holds:
$\quad\quad\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad\quad b_{n,k}\leq Ce^{m\alpha_{n}}\quad\quad\forall n\in\mathbb{N}.$ (4.1)
@@
The sufficiency part comes from Proposition 4.1. Assume $\theta\in(\Lambda_{\infty}(\alpha))^{\prime}$. Then there exists some $m_{1}\in\mathbb{N}$ and $C_{1}>0$ such that
$\quad\quad\quad\quad\quad\quad\quad\quad|\theta_{n-1}|\leq C_{1}e^{m_{1}\alpha_{n}}\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\forall n\in\mathbb{N}.$
From 4.1, for every $k\in\mathbb{N}$, there exist $m_{2}\in\mathbb{N}$ and $C_{2}>0$ such that
$\begin{split}\|\widecheck{T}_{\theta}e_{n}\|_{k}&=\sup_{1\leq j\leq n}|\theta_{n-j}|b_{j,k}\\ &\leq C_{1}\sup_{1\leq j\leq n}e^{m_{1}\alpha_{n-j+1}}b_{j,k}\\ &\leq C_{2}\sup_{1\leq j\leq n}e^{m_{1}\alpha_{n-j+1}}e^{m_{2}\alpha_{j}}\\ &\leq C_{2}\sup_{1\leq j\leq n}e^{m_{1}\alpha_{n-j+1}}e^{m_{2}\alpha_{n}}\\ &\leq C_{2}e^{m_{1}\alpha_{n}}e^{m_{2}\alpha_{n}}\\ &=C_{2}\|e_{n}\|_{m_{1}+m_{2}}.\end{split}$
Therefore, for all $k\in\mathbb{N}$, there exist $m=m_{1}+m_{2}\in\mathbb{N}$ such that
$\sup_{n}\frac{\|\widecheck{T}_{\theta}e_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty,$
that is, $\widecheck{T}_{\theta}:\Lambda_{\infty}(\alpha)\to K(b_{n,k})$ is continuous. ∎

### v3

@@ 3. Operators Defined by A Lower Triangular Toeplitz Matrix
If $\widehat{T}_{\theta}$ is compact, then $\widehat{T}_{\theta}$ is continuous and $\theta\in\Lambda_{1}(\beta)$ from Proposition 3.1. Proposition 3.2 says that the condition (3.2) holds. For the other direction, assume that $\theta\in\Lambda_{1}(\beta)$ and the condition (3.2) holds. By the same calculation of the proof of Theorem 3.3, we can write that there exists an $m\in\mathbb{N}$ such that for all $k\in\mathbb{N}$
@@
$\hskip 14.45377pt\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad e^{-\frac{1}{k}\beta_{n}}\leq Ca_{n,S(k)}\hskip 57.81621pt\forall n\in\mathbb{N}.$ (3.3)
@@
is $\tilde{S}$-tame, where $\tilde{S}(k)=S(2k)$ for every $k\in\mathbb{N}$.
@@
Let $\theta\in\Lambda_{1}(\beta)$. By using (3.3), for every $k\in\mathbb{N}$, there exist $C_{1},C_{2}>0$ such that
$\begin{split}\|\widehat{T}_{\theta}e_{n}\|_{k}&=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{2}{2k}\beta_{j}}=\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{2k}\beta_{j}}e^{-\frac{1}{2k}\beta_{j}}\\ &\leq\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{2k}\beta_{j}}\bigg{)}e^{-\frac{1}{2k}\beta_{n}}\leq C_{1}\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{2k}\beta_{j}}\bigg{)}a_{n,S(2k)}\\ &\leq C_{1}\bigg{(}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{-\frac{1}{2k}\beta_{j-n+1}}\bigg{)}\|e_{n}\|_{S(2k)}\leq C_{1}\|\theta\|_{2k}\|e_{n}\|_{\tilde{S}(k)}\leq C_{2}\|e_{n}\|_{\tilde{S}(k)}.\end{split}$
@@
$\|\widehat{T}_{\theta}e_{n}\|_{k}\leq C\|e_{n}\|_{\tilde{S}(k)},$
that is, the family $\mathcal{\widehat{A}}_{1}$ is $\tilde{S}$-tame. ∎
Now we turn our attention to the power series space of infinite type. For the converse of Proposition 3.1, we need the stability condition on the sequence $\beta$ when $K(b_{n,k})$ is power series space of infinite type $\Lambda_{\infty}(\beta)$. A sequence $\beta$ is called stable if
$\sup_{n\in\mathbb{N}}{\beta_{2n}\over\beta_{n}}<\infty.$
@@
Let $\beta=\left(\beta_{n}\right)_{n\in\mathbb{N}}$ be a stable sequence. $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{\infty}(\beta)$ is well-defined and continuous if and only if $\theta\in\Lambda_{\infty}(\beta)$ and the following condition holds:
$\hskip 14.45377pt\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad e^{k\beta_{n}}\leq Ca_{n,m}\hskip 50.58878pt\forall n\in\mathbb{N}.$ (3.4)
@@
Let assume $\theta$ be a sequence and $s_{0}=\min\{t:\theta_{t}\neq 0\}.$ If the operator $\widehat{T}_{\theta}$ is continuous, Proposition 3.1 gives us that $\theta\in\Lambda_{\infty}(\beta)$ and for every $k\in\mathbb{N}$, there exist $m\in\mathbb{N}$, $C>0$ such that
$e^{k\beta_{n+s_{0}}}\leq Ca_{n,m}$
and then we have
$e^{k\beta_{n}}\leq Ca_{n,m}$
for every $n\in\mathbb{N}$. This says that (3.4) holds since $\beta$ is an increasing sequence. For the converse, we assume that $\beta$ is stable. Then there exists an $M>1$ such that
$\hskip 119.24506pt\beta_{2n}\leq M\beta_{n}\hskip 126.47249pt\forall n\in\mathbb{N}.$
Since $\beta$ is increasing we have the following: if $j=2t\geq 2n$,
$\beta_{j}=\beta_{2t}\leq M\beta_{t}\leq M\beta_{2t-n}=M\beta_{j-n}\leq M\beta_{j-n+1}$
and if $j=2t+1\geq 2n$, we have $t+1\geq n$ and
$\beta_{j}=\beta_{2t+1}\leq\beta_{2t+2}\leq M\beta_{t+1}\leq M\beta_{2t+2-n}=M\beta_{j-n+1}.$
Therefore, we have
$\beta_{j}\leq M\beta_{j-n+1}$
for all $j\geq 2n$. Now let us assume that $\theta\in\Lambda_{\infty}(\beta)$ and the condition (3.4) holds. Then, for every $k\in\mathbb{N}$ there exist $m\in\mathbb{N}$, $C_{1},C_{2}>0$ such that
$\begin{split}\|\widehat{T}_{\theta}e_{n}\|_{k}&=\sum^{2n-1}_{j=n}|\theta_{j-n}|e^{k\beta_{j}}+\sum^{\infty}_{j=2n}|\theta_{j-n}|e^{k\beta_{j}}\\ &\leq\sum^{2n-1}_{j=n}|\theta_{j-n}|e^{k\beta_{2n}}+\sum^{\infty}_{j=2n}|\theta_{j-n}|e^{Mk\beta_{j-n+1}}e^{k\beta_{j}-Mk\beta_{j-n+1}}\\ &\leq e^{Mk\beta_{n}}\sum^{2n-1}_{j=n}|\theta_{j-n}|+\sum^{\infty}_{j=2n}|\theta_{j-n}|e^{Mk\beta_{j-n+1}}\leq e^{Mk\beta_{n}}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{Mk\beta_{j-n+1}}\\ &\leq C_{1}a_{n,m}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{Mk\beta_{j-n+1}}\leq C_{1}\|\theta\|_{Mk}\|e_{n}\|_{m}\leq C_{2}\|e_{n}\|_{m}.\end{split}$
holds for every $n\in\mathbb{N}$. We employed that $\beta_{j}\leq M\beta_{j-n+1}$ for every $j\geq 2n$ in the first line. Therefore, $\widehat{T}_{\theta}e_{n}\in\Lambda_{\infty}(\beta)$ for every $n\in\mathbb{N}$ and for all $k\in\mathbb{N}$ there exist $m\in\mathbb{N}$ such that
@@
that is, $\widehat{T}_{\theta}$ is well-defined and continuous by Proposition 2.2. ∎
@@
Let $\beta=\left(\beta_{n}\right)_{n\in\mathbb{N}}$ be a stable sequence. $\widehat{T}_{\theta}:K(a_{n,k})\to\Lambda_{\infty}(\beta)$ is compact if and only if $\theta\in\Lambda_{\infty}(\beta)$ and the following condition holds:
$\hskip 21.68121pt\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad e^{k\beta_{n}}\leq Ca_{n,m}\hskip 21.68121pt\forall n\geq k.$ (3.5)
@@
Let assume $\theta$ be a sequence and $s_{0}=\min\{t:\theta_{t}\neq 0\}.$ If $\widehat{T}_{\theta}$ is compact, then $\theta\in\Lambda_{1}(\beta)$ from Proposition 3.1. Proposition 3.2 says that there exists a $m\in N$ such that for all $k\in\mathbb{N}$ there exists a $C>0$ such that
$e^{k\beta_{n+s_{0}}}\leq Ca_{n,m}$
for every $n\in\mathbb{N}$. This says that the condition (3.5) is satisfied since $\beta$ is an increasing sequence.
For the converse, assume that $\theta\in\Lambda_{\infty}(\beta)$ and (3.5) holds. By the same calculation of the proof of Theorem 3.6, we can write that there exists $m\in\mathbb{N}$ such that for all $k\in\mathbb{N}$
$\sup_{n\geq k}\frac{\|\widehat{T}_{\theta}e_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty$
and then
@@
Let $\beta=\left(\beta_{n}\right)_{n\in\mathbb{N}}$ be a non-negative increasing stable sequence with $\lim_{n\rightarrow\infty}\beta_{n}=\infty$. Assume that there exists a non-decreasing function $S:\mathbb{N}\to\mathbb{N}$ such that
$\hskip 14.45377pt\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad\quad e^{k\beta_{n}}\leq Ca_{n,S(k)}\hskip 57.81621pt\forall n\in\mathbb{N}.$ (3.6)
@@
is $\tilde{S}$-tame, where $\tilde{S}(k)=S(Mk)$ for every $k\in\mathbb{N}$ and here $M=\sup_{n\in\mathbb{N}}\frac{\alpha_{2n}}{\alpha_{n}}$.
@@
The proof is similar to the proof of Theorem 3.7. Let $\theta\in\Lambda_{\infty}(\beta)$. In the proof of Theorem 3.6 we showed that for all $k,n\in\mathbb{N}$
$\|\widehat{T}_{\theta}e_{n}\|_{k}\leq e^{Mk\beta_{n}}\sum^{\infty}_{j=n}|\theta_{j-n}|e^{Mk\beta_{j-n+1}}\leq e^{Mk\beta_{n}}\|\theta\|_{Mk}.$
Using the condition (3.6), we can write that for all $k\in\mathbb{N}$ there exists a $C>0$ such that
$\|\widehat{T}_{\theta}e_{n}\|_{k}\leq e^{Mk\beta_{n}}\|\theta\|_{Mk}\leq Ca_{n,S(Mk)}=Ca_{n,\tilde{S}(k)}.$
@@
$\hskip 72.26999pt\forall k\quad\exists C>0\hskip 50.58878pt\|\widehat{T}_{\theta}e_{n}\|_{k}\leq C\|e_{n}\|_{\tilde{S}(k)}\hskip 72.26999pt$
@@ 4. Operators Defined by an Upper Triangular Toeplitz Matrix
We want to define a linear operator $\widecheck{T}_{\theta}:K(a_{n,k})\to K(b_{n,k})$ by taking $\widecheck{T}_{\theta}e_{n}$ as the nth column of the above matrix, that is,
@@
In fact, we can not guarantee that the operator $\widecheck{T}_{\theta}$ is correctly defined between two general Köthe spaces $K(a_{n,k})$ and $K(b_{n,k})$, because we do not know if the series $\sum^{\infty}_{n=1}x_{n}\widecheck{T}_{\theta}e_{n}$ converges in $K(b_{n,k})$ for every $x\in K(a_{n,k})$ in this general case. Below, we will examine the cases where this operator can be correctly defined, and in these cases, we will analyze the continuity and compactness of this operator. Similiar to the previous section, we will assume that the sequences $\theta$ satisfy the following condition
$\forall n\in\mathbb{N}\hskip 14.45377pt\exists s>n\hskip 28.90755pt\theta_{s}\neq 0.$
Otherwise, $\theta$ produces a finite rank operator, which is obviously continuous and compact.
In [10, Theorem 2.2], you can find the same idea for Proposition 4.1 and a generalized form of Theorem 4.3 for $G_{\infty}$-spaces.
@@
Let $K(a_{n,k})$, $K(b_{n,k})$ be Köthe spaces and assume that $\theta$ be a sequence and $s_{0}=\min\{t:\theta_{t}\neq 0\}$. If $\widecheck{T}_{\theta}:K(a_{n,k})\to K(b_{n,k})$ is continuous, then $\theta\in(K(a_{n,k}))^{\prime}$ and the following holds:
$\hskip 18.06749pt\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad\quad b_{n-s_{0},k}\leq Ca_{n,m}\hskip 46.97505pt\forall n>s_{0}.$
@@
Since $(b_{n,k})$ is a Köthe matrix, there exists a $k_{0}\in\mathbb{N}$ satisfying $b_{1,k_{0}}\neq 0$. Let us take $C_{2}={C_{1}\over b_{1,k_{0}}}$. Then we can write
@@
and this says that $\theta\in(K(a_{n,k}))^{\prime}$. Further, we have that for every $k\in\mathbb{N}$, there exist $m\in\mathbb{N}$ and $C_{1}>0$ so that
$|\theta_{s_{0}}|b_{n-s_{0}}\leq C_{1}a_{n,m}$
and
$b_{n-s_{0}}\leq\frac{C_{1}}{\theta_{s_{0}}}a_{n,m}.$
hold for all $n>s_{0}$. By choosing $C=\frac{C_{1}}{|\theta_{s_{0}}|}$, we have
$\hskip 108.405ptb_{n-s_{0},k}\leq C_{3}a_{n,m}\hskip 108.405pt\forall n>s_{0}.$
@@
For compact $\widecheck{T}_{\theta}:K(a_{n,k})\to K(b_{n,k})$ operators, the relationship between Köthe matrices $(a_{n,k})_{n,k\in\mathbb{N}}$ and $(b_{n,k})_{n,k\in\mathbb{N}}$ is as follows:
Proposition 4.2.
Let $K(a_{n,k})$ and $K(b_{n,k})$ be Köthe spaces such that $K(b_{n,k})$ is Montel and assume that $\theta\in K(b_{n,k})$ and $s_{0}=\min\{t:\theta_{t}\neq 0\}$. If $\widecheck{T}_{\theta}:K(a_{n,k})\to K(b_{n,k})$ is a compact operator, then the following holds
$\hskip 25.29494pt\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad\quad b_{n-s_{0},k}\leq Ca_{n,m}\hskip 36.135pt\forall n>s_{0}.$
Proof.
The proof is similar to the proof of Proposition 4.1. ∎
@@
Let $\Lambda_{\infty}(\alpha)$ be a nuclear power series space of infinite type, $K(b_{n,k})$ be a Köthe space and assume that $\theta$ be a sequence and $s_{0}=\min\{t:\theta_{t}\neq 0\}$. The operator $\widecheck{T}_{\theta}:\Lambda_{\infty}(\alpha)\to K(b_{n,k})$ is well-defined and continuous if and only if $\theta\in(\Lambda_{\infty}(\alpha))^{\prime}$ and the following condition holds:
$\hskip 14.45377pt\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad b_{n-s_{0},k}\leq Ce^{m\alpha_{n}}\hskip 28.90755pt\forall n>s_{0}.$ (4.1)
@@
If the operator $\widecheck{T}_{\theta}$ is continuous, Proposition 4.1 gives us that $\theta\in(\Lambda_{\infty}(\alpha))^{{}^{\prime}}$ and (4.1) holds. Let us assume that $\theta\in(\Lambda_{\infty}(\alpha))^{\prime}$ and (4.1) holds. Then there exists some $m_{1}\in\mathbb{N}$ and $C_{1}>0$ such that
$\hskip 108.405pt|\theta_{n-1}|\leq C_{1}e^{m_{1}\alpha_{n}}\hskip 112.0187pt\forall n\in\mathbb{N}.$
From the condition (4.1), for every $k\in\mathbb{N}$, there exist $m_{2}\in\mathbb{N}$ and $C_{2}>0$ such that
$\begin{split}\|\widecheck{T}_{\theta}e_{n}\|_{k}&=\sum^{n}_{j=1}|\theta_{n-j}|b_{j,k}=\sum^{n-s_{0}}_{j=1}|\theta_{n-j}|b_{j,k}\leq C_{1}\sum^{n-s_{0}}_{j=1}e^{m_{1}\alpha_{n-j+1}}b_{j,k}\\ &\leq C_{2}\sum^{n-s_{0}}_{j=1}e^{m_{1}\alpha_{n-j+1}}e^{m_{2}\alpha_{j+s_{0}}}\leq C_{2}e^{m_{1}\alpha_{n}}\sum^{n-s_{0}}_{j=1}e^{m_{2}\alpha_{j+s_{0}}}\\ &=C_{2}e^{m_{1}\alpha_{n}}\sum^{n}_{j=1+s_{0}}e^{m_{2}\alpha_{j}}.\end{split}$
On the other hand, $\Lambda_{\infty}(\alpha)$ is nuclear then $\sum^{\infty}_{n=1}e^{-m_{3}\alpha_{n}}$ is convergent for some $m_{3}\in\mathbb{N}$ and then for some $D>0$ we have
$\sum^{n}_{j=1}e^{m_{2}\alpha_{j}}\leq De^{(m_{2}+m_{3})\alpha_{n}}$
since
$\sum^{n}_{j=1}e^{m_{2}\alpha_{j}}e^{-(m_{2}+m_{3})\alpha_{n}}\leq\sum^{n}_{j=1}e^{m_{2}\alpha_{j}}e^{-(m_{2}+m_{3})\alpha_{j}}\leq\sum^{\infty}_{j=1}e^{-m_{3}\alpha_{j}}=D<+\infty.$
Then we can write
$\begin{split}\|\widecheck{T}_{\theta}e_{n}\|_{k}&\leq C_{2}e^{m_{1}\alpha_{n}}\sum^{n}_{j=1+s_{0}}e^{m_{2}\alpha_{j}}\leq C_{2}e^{m_{1}\alpha_{n}}\sum^{n}_{j=1}e^{m_{2}\alpha_{j}}\\ &\leq DC_{2}e^{m_{1}\alpha_{n}}e^{(m_{2}+m_{3})\alpha_{n}}=DC_{2}\|e_{n}\|_{m_{1}+m_{2}+m_{3}}.\end{split}$
Therefore, $\widecheck{T}_{\theta}e_{n}\in K(b_{n,k})$ for every $n\in\mathbb{N}$ and for all $k\in\mathbb{N}$, there exist $m=m_{1}+m_{2}+m_{3}\in\mathbb{N}$ such that
that is, $\widecheck{T}_{\theta}$ is well defined and continuous by Proposition 2.2. ∎